OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#12490 — mtl-1-6k
Scheduled Maintenance Report for Network & Infrastructure
Completed
Link flap between BHS and MTL, we are going to replace the optics on BHS side.

Update(s):

Date: 2015-02-19 15:53:15 UTC
The flap is back again.

We've found a faulty connection on a patch. The flap is patched. We will continue to monitor it.

Date: 2015-02-09 12:34:33 UTC
We've replaced the jumper and the connection is back up. We will monitor it.

Date: 2015-02-05 09:28:13 UTC
The optic cable has been replaced in MTL, but the connection is still down. We'll change the jumpers asap.

Date: 2015-02-04 15:15:22 UTC
An optic cable has been replaced at MTL. If this doesn't fix the issue, we will change all the jumpers in this area.

Date: 2015-02-02 19:53:59 UTC
Again, flaps appeared. We will replace the entire optical path on MTL, the same action had no result on BHS.

Date: 2015-01-28 09:28:04 UTC
We continue to monitor the ports in question.

Date: 2015-01-27 14:09:36 UTC
The link flap is again present , we will work on the
mtl site to clean the lens ans optics .

Date: 2015-01-27 13:41:21 UTC
We are going to proceed with replacement of the concerned optics.
We are monitoring the CRC and flaps on the interface in order to close this task.
Posted Jan 27, 2015 - 13:40 UTC